A man rows directly across a river in time t second and rows an equal distance down the stream in T second. Find the ratio of man's speed in still water to the speed of river water

Grade:12th pass

3 Answers

Mohit Chhabra
8 years ago
let the man rows a distance of d meters
let velocity of man is v in still water
let velocity of river is V
 
now time taken by man to row in still water is t, which is equal to
t=d/v {eq 1}
now time taken by man to row the same distance while going downstream is T, which is equal to
T=d/(v+V) {eq 2}  {velocity of man and river adds while going downstream}
divide numerator and denominator in eq 2 by v
1+(V/v) = t/T
 
v/V = T/(t-T)
pls tell is this ansewr helpful..
Abhinav
8 years ago
let speed of man be V
and speed of river flow U.

thanV=\frac{d}{t }

and

V+U=\frac{d}{T}

dividing both equations and rearranging


\frac{U}{V}=\frac{t}{T}-1






Arati Sinha
8 years ago
Sorry, mohit and abhinav but the answer is( t2+T2)/(t2-T2). So the process suggested is not helpful. I hope after knowing the answer maybe you can suggest me the correct process.....